# Generative Uncertainty in Diffusion Models

### v1

@@ A.1 Qualitative Comparison with BayesDiff
To further highlight the differences between our generative uncertainty and BayesDiff [Kou et al., 2024], we present samples with the highest and lowest uncertainty according to BayesDiff in Figure 5. These samples are drawn from the same set of 12K ImageNet ‘unconditional’ images generated using the UViT model [Bao et al., 2023] as in Figure 2. Notably, BayesDiff’s uncertainty score appears highly sensitive to background pixels—images with high uncertainty tend to have cluttered backgrounds, while those with low uncertainty typically feature clear backgrounds. Furthermore, as reflected in BayesDiff’s poor performance in terms of FID and precision (see Table 1), some low-uncertainty examples exhibit noticeable artefacts, whereas certain high-uncertainty samples are of rather high-quality. For example, the image of a dog in the bottom-right corner of the high-uncertainty grid in Figure 5 looks quite good despite being assigned (very) high uncertainty.
@@
A.2 Ablation on Semantic Likelihood
To highlight the importance of using a semantic likelihood (Section 3.3) when leveraging uncertainty to detect low-quality generations, we conduct an ablation study in which we replace it with a standard Gaussian likelihood applied directly in pixel space (Eq. 6). Figure 7 presents the highest and lowest uncertainty images according to this ‘pixel-space’ generative uncertainty. Notably, pixel-space uncertainty is overly sensitive to background pixels, mirroring the issue observed in BayesDiff (see Appendix A.1). This highlights the necessity of using semantic likelihood to obtain uncertainty estimates that are truly informative about the visual quality of generated samples.
@@ A.2 Ablation on Semantic Likelihood
A.3 Pixel-Wise Uncertainty
@@
Although pixel-wise uncertainty received significant attention in past work [Kou et al., 2024, Chan et al., 2024, De Vita and Belagiannis, 2025], there is currently no principled method for evaluating its quality. Most existing approaches rely on qualitative inspection, visualizing pixel-wise uncertainty for a few generated samples (as we do in Figure 8). This further motivates our focus on sample-wise uncertainty estimates, where more rigorous evaluation frameworks—such as improvements in FID and precision on a set of filtered images (see Table 1)—enable more meaningful comparisons between different approaches.
@@
A.4 Comparison with Likelihood
We compare our generative uncertainty filtering criterion with a likelihood selection approach on the 12K images generated by ADM trained on ImageNet 128x128. In the same way as in our other comparisons, we retain the 10K generated images with highest likelihood. We utilize the implementation in Dhariwal and Nichol [2021] to compute the bits-per-dimension of each sample (one-to-one with likelihood). The 25 samples with lowest and highest likelihood are shown in Figure 9. Visually, the likelihood objective heavily prefers simple images with clean backgrounds and not necessarily image quality. Note that this is consistent with other works that have reported likelihood to be an inconsistent identifier of image quality [Theis et al., 2016, Theis, 2024]. Quantitative results for image quality were consistent with our qualitative observations. The FID, precision, and recall for the best 10K images according to bits-per-dimension were $11.86\pm 0.0026$, $58.23\pm 0.02160$, and $70.45\pm 0.0237$ over three runs. By point estimate, all three metrics are worse or indistinguishable from the Random baseline ($11.31\pm 0.07,58.90\pm 0.36,70.68\pm 0.38$). Results for all image selection methods can be found in Table 1.
@@ A.4 Comparison with Likelihood
A.5 Comparison with Realism & Rarity
@@
In contrast, our generative uncertainty exhibits a weaker correlation ($<0.4$) with both realism and rarity scores. We attribute this to the fact that our uncertainty primarily reflects the limited training data used in training diffusion models (i.e., epistemic uncertainty, see Section 3.4), rather than the distance to a reference dataset, as is the case for realism and rarity scores.
@@
Table 2: Image generation results for 10K filtered samples (out of 12K) based on combined metrics. Combining our generative uncertainty outperforms combining realism and recall in terms of FID. We report mean values along with standard deviation over 3 runs with different random seeds.
@@
A.6 Class-averaged generative uncertainty
To better understand the drop in sample diversity (recall) when using our generative uncertainty to filter low-quality samples in Table 1, we analyze the distribution of average entropy per conditioning class. Specifically, for each of the 12K generated images, we randomly sample a conditioning class to mimic unconditional generation. As a result, all 1,000 ImageNet classes are represented among the 12K generated samples. Next, we compute our generative uncertainty for each sample and then average the uncertainties within each class. A plot of class-averaged uncertainties is shown in Figure 11. Since class-averaged uncertainties exhibit considerable variance, the class distribution in the 10K filtered samples deviates somewhat from that of the original 12K images, thereby explaining the reduction in diversity (recall).
@@ A.6 Class-averaged generative uncertainty
A.7 Flow Matching
To demonstrate that our generative uncertainty framework (Section 3) extends beyond diffusion models, we apply it here to the recently popularized flow matching approach [Lipman et al., 2022, Liu et al., 2022, Albergo et al., 2023]. Specifically, we consider a latent flow matching formulation [Dao et al., 2023] with a DiT backbone [Peebles and Xie, 2023]. For sampling, we employ a fifth-order Runge-Kutta ODE solver (dopri5). In Figure 12, we illustrate the samples with the highest and lowest generative uncertainty among 12K generated samples. On a filtered set of 10K images, our generative uncertainty framework achieves an FID of $10.48$ and a precision of $64.71$, significantly outperforming a random baseline, which yields an FID of $11.80$ and a precision of $61.04$.
Figure 12: Images with the highest (left) and the lowest (right) generative uncertainty (Eq. 3.3) among 12K generations using a latent flow matching model [Dao et al., 2023]. Generative uncertainty correlates with visual quality, as high-uncertainty samples exhibit numerous artefacts, whereas low-uncertainty samples resemble canonical images of their respective conditioning class.
Appendix B Implementation Details
All our experiments can be conducted on a single $A100$ GPU, including the fitting of the Laplace posterior (Section 3.2). Code for reproducing our experiments is publicly available at https://github.com/metodj/DIFF-UQ.
@@ Appendix B Implementation Details
When fitting a last-layer Laplace approximation (Section 3.2), we closely follow the implementation from BayesDiff [Kou et al., 2024]. Specifically, we use the empirical Fisher approximation with a diagonal factorization for Hessian computation. The prior precision parameter and observation noise are fixed at $\gamma=1$ and $\sigma=1$, respectively. For Hessian computation, we utilize $1\%$ of the training data for ImageNet 128×128 and $2\%$ for ImageNet 256×256. Further details about the last layer of each diffusion model are provided in Table 3, where we observe that fewer than $1\%$ of the parameters receive a ‘Bayesian treatment’. We utilize laplace [^4] https://github.com/aleximmer/Laplace library in our implementation.
@@
For our main experiment in Section 4.1, we generate 12K images using the pretrained ADM model [Dhariwal and Nichol, 2021] for ImageNet 128×128 and the UViT model [Bao et al., 2023] for ImageNet 256×256. Following BayesDiff [Kou et al., 2024], we use a DDIM sampler [Song et al., 2020a] for the ADM model and a DPM-2 sampler [Lu et al., 2022] for the UViT model, both with $T=50$ sampling steps.

### v2

@@ C.1 Qualitative Comparison with BayesDiff
To further highlight the differences between our generative uncertainty and BayesDiff [Kou et al., 2024], we present samples with the highest and lowest uncertainty according to BayesDiff in Figure 8. These samples are drawn from the same set of 12K ImageNet images generated using the UViT model [Bao et al., 2023] as in Figure 4. Notably, BayesDiff’s uncertainty score appears highly sensitive to background pixels—images with high uncertainty tend to have cluttered backgrounds, while those with low uncertainty typically feature clear backgrounds. Furthermore, as reflected in BayesDiff’s poor performance in terms of FID and precision (see Figures 3&7), some low-uncertainty examples exhibit noticeable artefacts, whereas certain high-uncertainty samples are of rather high-quality. For example, the image of a dog in the bottom-right corner of the high-uncertainty grid in Figure 8 looks quite good despite being assigned (very) high uncertainty.
@@
C.2 Ablation on Semantic Likelihood
To highlight the importance of using a semantic likelihood (Section 3.3) when leveraging uncertainty to detect low-quality generations, we conduct an ablation study in which we replace it with a standard Gaussian likelihood applied directly in pixel space (Eq. 6). Figure 10 presents the highest and lowest uncertainty images according to this ‘pixel-space’ generative uncertainty. Notably, pixel-space uncertainty is overly sensitive to background pixels, mirroring the issue observed in BayesDiff (see Appendix C.1). This highlights the necessity of using semantic likelihood to obtain uncertainty estimates that are truly informative about the visual quality of generated samples.
@@ C.2 Ablation on Semantic Likelihood
C.3 Pixel-Wise Uncertainty
@@
Although pixel-wise uncertainty received significant attention in past work [Kou et al., 2024, Chan et al., 2024, De Vita and Belagiannis, 2025], there is currently no principled method for evaluating its quality. Most existing approaches rely on qualitative inspection, visualizing pixel-wise uncertainty for a few generated samples (as we do in Figure 11). This further motivates our focus on sample-wise uncertainty estimates, where more rigorous evaluation frameworks—such as improvements in FID and precision on a set of filtered images—enable more meaningful comparisons between different approaches.
@@
C.4 Comparison with Likelihood
We compare our generative uncertainty filtering criterion with a likelihood selection approach on the 12K images generated by ADM trained on ImageNet 128x128. Here retain the $n=10\textrm{K}$ generated images with highest likelihood. We utilize the implementation in Dhariwal and Nichol [2021] to compute the bits-per-dimension of each sample (one-to-one with likelihood). The 25 samples with lowest and highest likelihood are shown in Figure 12. Visually, the likelihood objective heavily prefers simple images with clean backgrounds and not necessarily image quality. Note that this is consistent with other works that have reported likelihood to be an inconsistent identifier of image quality [Theis et al., 2016, Theis, 2024]. Quantitative results for image quality were consistent with our qualitative observations. The FID, precision, and recall for the best 10K images according to bits-per-dimension were $11.86\pm 0.0026$, $58.23\pm 0.02160$, and $70.45\pm 0.0237$ over three runs. By point estimate, all three metrics are worse or indistinguishable from the Random baseline ($11.31\pm 0.07,58.90\pm 0.36,70.68\pm 0.38$).
@@ C.4 Comparison with Likelihood
C.5 Comparison with Realism & Rarity
Figure 13: Image generation results for $n\in\{6\textrm{K},7\textrm{K},\ldots,11\textrm{K}\}$ filtered samples (out of 12K) for ADM diffusion model [Dhariwal and Nichol, 2021]. Our generative uncertainty performs on par with realism [Kynkäänniemi et al., 2019] and rarity scores [Han et al., 2023] in terms of FID (left), while exhibiting a weaker precision-recall trade-off (middle and right). We report mean values along with standard deviations over 5 runs with different random seeds.
Figure 14: Image generation results for $n\in\{6\textrm{K},7\textrm{K},\ldots,11\textrm{K}\}$ filtered samples (out of 12K) for UViT diffusion model [Bao et al., 2023]. Our generative uncertainty performs on par with realism [Kynkäänniemi et al., 2019] and rarity scores [Han et al., 2023] in terms of FID (left), while exhibiting a weaker precision-recall trade-off (middle and right). We report mean values along with standard deviations over 5 runs with different random seeds.
@@ C.5 Comparison with Realism & Rarity
In contrast, our generative uncertainty exhibits a weaker correlation ($<0.4$) with both realism and rarity scores. We attribute this to the fact that our uncertainty primarily reflects the limited training data used in training diffusion models (i.e., epistemic uncertainty), rather than the distance to a reference dataset, as is the case for realism and rarity scores.
@@
[…]
@@
C.6 Class-averaged generative uncertainty
@@
To better understand the drop in sample diversity (recall) when using our generative uncertainty to filter low-quality samples in Figures 3&7, we analyze the distribution of average entropy per conditioning class. Specifically, for each of the 12K generated images, we randomly sample a conditioning class to mimic unconditional generation. As a result, all 1,000 ImageNet classes are represented among the 12K generated samples. Next, we compute our generative uncertainty for each sample and then average the uncertainties within each class. A plot of class-averaged uncertainties is shown in Figure 16. Since class-averaged uncertainties exhibit considerable variance, the class distribution in the 10K filtered samples deviates somewhat from that of the original 12K images, thereby explaining the reduction in diversity (recall).
@@
C.7 Flow Matching
To demonstrate that our generative uncertainty framework (Section 3) extends beyond diffusion models, we apply it here to the recently popularized flow matching approach [Lipman et al., 2023, Liu et al., 2023, Albergo et al., 2023]. Specifically, we consider a latent flow matching formulation [Dao et al., 2023] with a DiT backbone [Peebles and Xie, 2023]. For sampling, we employ a fifth-order Runge-Kutta ODE solver (dopri5). In Figure 17, we illustrate the samples with the highest and lowest generative uncertainty among 12K generated samples. On a filtered set of 10K images, our generative uncertainty framework achieves an FID of $10.48$ and a precision of $64.71$, significantly outperforming a random baseline, which yields an FID of $11.80$ and a precision of $61.04$.
Figure 17: Images with the highest (left) and the lowest (right) generative uncertainty among 12K generations using a latent flow matching model [Dao et al., 2023]. Uncertainty correlates with visual quality, as high-uncertainty samples exhibit numerous artefacts, whereas low-uncertainty samples resemble canonical images of their respective conditioning class.
Appendix D Implementation Details
All our experiments can be conducted on a single $A100$ GPU, including the fitting of the Laplace posterior (Section 3.2). Our code is publicly available at https://github.com/metodj/DIFF-UQ.
@@ Appendix D Implementation Details
When fitting a last-layer Laplace approximation (Section 3.2), we closely follow the implementation from BayesDiff [Kou et al., 2024]. Specifically, we use the empirical Fisher approximation with a diagonal factorization for Hessian computation. As the prior, we adopt a simple isotropic Gaussian distribution, $p(\theta)=\mathcal{N}(0,\gamma^{-1}I)$. The prior precision parameter and observation noise are fixed at $\gamma=1$ and $\sigma=1$, respectively. We report ablations for both parameters in Tables 4 and 4, finding that neither has a significant impact on the results. For Hessian computation, we utilize $1\%$ of the training data for ImageNet 128×128 and $2\%$ for ImageNet 256×256. Further details about the last layer of each diffusion model are provided in Table 2, where we observe that fewer than $1\%$ of the parameters receive a ‘Bayesian treatment’. We utilize laplace [^4] https://github.com/aleximmer/Laplace library in our implementation.
@@
For our main experiment in Section 4.2, we generate 12K images using the pretrained ADM model [Dhariwal and Nichol, 2021] for ImageNet 128×128 and the UViT model [Bao et al., 2023] for ImageNet 256×256. Following BayesDiff [Kou et al., 2024], we use a DDIM sampler [Song et al., 2021a] for the ADM model and a DPM-2 sampler [Lu et al., 2022] for the UViT model, both with $T=50$ sampling steps.
@@
Table 3: Ablation on prior precision parameter $\gamma$. Results for ADM model on ImageNet 128x128 dataset based on $n=10\textrm{K}$ filtered images (out of 12K).
$\gamma$ FID ()($\downarrow$)( ↓ ) Precision ()($\uparrow$)( ↑ ) Recall ()($\uparrow$)( ↑ )
$1.0$ $10.04\pm 0.14$ $61.28\pm 0.23$ $69.55\pm 0.49$
$0.01$ $10.04\pm 0.12$ $61.14\pm 0.25$ $69.59\pm 0.52$
$0.1$ $10.05\pm 0.09$ $61.19\pm 0.21$ $69.75\pm 0.45$
$10$ $10.01\pm 0.15$ $60.95\pm 0.28$ $69.71\pm 0.54$
$100$ $10.06\pm 0.11$ $61.12\pm 0.26$ $69.62\pm 0.50$
Table 4: Ablation on likelihood noise $\sigma^{2}$ parameter. Results for ADM model on ImageNet 128x128 dataset based on $n=10\textrm{K}$ filtered images (out of 12K).
$\sigma^{2}$ FID ()($\downarrow$)( ↓ ) Precision ()($\uparrow$)( ↑ ) Recall ()($\uparrow$)( ↑ )
$0.1$ $10.30\pm 0.11$ $60.62\pm 0.31$ $69.99\pm 0.43$
$0.01$ $10.18\pm 0.04$ $61.18\pm 0.38$ $69.61\pm 0.59$
$0.001$ $10.04\pm 0.14$ $61.28\pm 0.23$ $69.55\pm 0.49$
$0.0001$ $10.01\pm 0.14$ $61.34\pm 0.23$ $69.50\pm 0.39$
$0.00001$ $10.10\pm 0.06$ $61.40\pm 0.28$ $69.53\pm 0.57$
Additional Acknowledgements and Disclaimers
Parts of this research were supported by the Intelligence Advanced Research Projects Activity (IARPA) via Department of Interior/ Interior Business Center (DOI/IBC) contract number 140D0423C0075. The U.S. Government is authorized to reproduce and distribute reprints for Governmental purposes notwithstanding any copyright annotation thereon. Disclaimer: The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies or endorsements, either expressed or implied, of IARPA, DOI/IBC, or the U.S. Government.